# Deciphering Hate: Identifying Hateful Memes and Their Targets

### v1

@@ Abstract
Internet memes have become a powerful means for individuals to express emotions, thoughts, and perspectives on social media. While often considered as a source of humor and entertainment, memes can also disseminate hateful content targeting individuals or communities. Most existing research focuses on the negative aspects of memes in high-resource languages, overlooking the distinctive challenges associated with low-resource languages like Bengali (also known as Bangla). Furthermore, while previous work on Bengali memes has focused on detecting hateful memes, there has been no work on detecting their targeted entities. To bridge this gap and facilitate research in this arena, we introduce a novel multimodal dataset for Bengali, BHM (Bengali Hateful Memes). The dataset consists of 7,148 memes with Bengali as well as code-mixed captions, tailored for two tasks: (i) detecting hateful memes, and (ii) detecting the social entities they target (i.e., Individual, Organization, Community, and Society). To solve these tasks, we propose DORA (Dual cO-attention fRAmework), a multimodal deep neural network that systematically extracts the significant modality features from the memes and jointly evaluates them with the modality-specific features to understand the context better. Our experiments show that DORA is generalizable on other low-resource hateful meme datasets and outperforms several state-of-the-art rivaling baselines.
@@ 1 Introduction
In recent years, social media has brought a distinct form of multimodal entity: memes, providing a means to express ideas and emotions. Memes are compositions of images coupled with concise text. While memes are often amusing, they can also spread hate by incorporating socio-political elements. These hateful memes pose a significant threat to social harmony as they have the potential to harm individuals or specific groups based on factors like political beliefs, sexual orientation, or religious affiliations. Despite the significant influence of memes, their multifaceted nature and concealed semantics make them very hard to analyze. The prevalence of highly toxic memes in recent times has led to a growing body of research into the negative aspects of memes, such as hate Kiela et al. (2020), offensiveness Shang et al. (2021), and harm Pramanick et al. (2021b). However, most works focused on the memes in high-resource languages while only a few studied the objectionable (i.e., hate, abuse) memes of low-resource languages Kumari et al. (2023). This is particularly true for Bengali.
@@
To bridge this research gap, we develop a novel Bengali meme dataset encompassing the targeted entities of hate. The captions in the dataset contain code-mixed (Bengali + English) and code-switched text (written Bengali dialects in English alphabets). This makes the dataset more distinctive and challenging compared to previous studies in the field. On the technical front, prior research on hateful meme detection Kiela et al. (2019); Pramanick et al. (2021a) revealed that off-the-shelf multimodal systems, which often perform well on a range of visual-linguistic tasks, struggle when applied to memes. Besides, most of the existing visio-linguistic models Radford et al. (2021); Li et al. (2019, 2022) are primarily trained on image-text pairs of English languages, thus limiting their capability on low-resource languages. Moreover, the existing state-of-the-art multimodal models Pramanick et al. (2021c); Lee et al. (2021) for hateful meme detection can not be replicated because several components of their architectures are not available in low-resource languages (i.e, Bengali). To tackle these issues, we develop a multimodal framework and our major contributions are as follows.
  • We develop a benchmark multimodal dataset comprising 7,148 Bengali memes. The dataset includes two sets of labels for (i) detecting hateful memes and (ii) identifying targeted entities (individuals, organizations, communities, and society). We also provide detailed annotation guidelines to facilitate resource creation in this domain for other low-resource languages.
@@ 2 Related Work
Hateful memes dataset: Over the past few years, several meme datasets have been developed in the context of hate speech and its various intensity levels such as offense, harm, abuse, and troll. Sabat et al. (2019) developed a hateful memes dataset containing 5,020 memes. Facebook AI Kiela et al. (2020) contributed to developing a hateful memes dataset consisting of around 10K synthetic memes labeled in hateful and not-hateful classes. Similarly, another large-scale dataset comprising 150K memes was introduced by Gomez et al. (2020) for hate speech detection. In another work, Suryawanshi et al. (2020) developed an offensive memes dataset comprising 743 memes collected during the event of the 2016 US presidential election. Pramanick et al. (2021a) built a dataset to detect harmful memes containing around 3.5K memes related to COVID-19.
@@
Some works have also been conducted concerning low-resource languages. Perifanos and Goutsos (2021) developed a dataset consisting of 4,004 memes for detecting hate speech in the Greek language. Kumari et al. (2023) developed an offensive memes dataset consisting of 7,417 Hindi memes. Recently, Das and Mukherjee (2023) introduced a dataset comprising 4,043 samples for detecting the Bengali abusive memes. Two prior studies focused on hateful meme detection in the Bengali language. Karim et al. (2022) introduced a synthetic hate speech dataset comprising around 4,500 Bengali memes. Similarly, Hossain et al. (2022) developed another Bengali hateful memes dataset having 4,158 memes labeled hateful or not-hateful. However, none of these datasets annotated the targets of hateful memes.
@@
Multimodal hateful meme detection: Over the years, various approaches have been employed for detecting hate speech using multimodal learning. Earlier researchers utilized conventional fusion (i.e., early and late) Suryawanshi et al. (2020); Gomez et al. (2020) for producing multimodal representation. Later, some researchers employed bilinear pooling (Chandra et al., 2021) while others developed transformer-based multimodal architectures such as MMBT Kiela et al. (2019), ViLBERT Lu et al. (2019), and Visual-BERT Li et al. (2019). Besides, some works attempted to use disentangled learning (Lee et al., 2021), incorporate additional caption (Zhou et al., 2021), and add external knowledge Pramanick et al. (2021c) to improve the hateful memes detection performance. Recently, Cao et al. (2022) applied prompting techniques for hateful meme detection in English.
@@
Differences with existing studies: Though a considerable amount of studies have been conducted on hateful meme detection, only a few studies have focused on Bengali. We point out several drawbacks in the existing research on Bengali. Firstly, the existing hateful memes datasets are small. They framed the task as a binary (hateful or not-hateful) classification problem, overlooking the social entities (i.e., individuals, society) that a hateful meme can target. Our dataset provides both levels of annotation. The richness of our dataset contributes to a comprehensive understanding of the dynamics of Bengali memes. Only one particular work Das and Mukherjee (2023) studied the targets of abusive memes in Bengali. Their work is completely different from ours as we are attempting to identify targeted social entities of hateful memes, which is more explicit than the term abuse. Secondly, the current research overlooked the memes containing cross-lingual captions, while a large number of internet memes are written in code-mixed and code-switched manner. Lastly, none of the works provided any model that can be generalized across low-resource datasets.
@@ 3 BHM: A New Benchmark Dataset
As per our exploration, there is currently no existing dataset in Bengali focused on capturing hateful memes that target specific entities. To fill this gap, we develop a novel multimodal hateful meme dataset. We followed the guidelines outlined by Hossain et al. (2022) and Kiela et al. (2020) to develop the dataset. This section will provide a detailed discussion of the dataset development steps, encompassing the data collection and annotation process, along with relevant statistics.
@@ 3.1 Data Collection and Sampling
We collected memes from various online platforms, including Facebook, Instagram, Pinterest, and Blogs. We used keywords such as "Bengali Troll Memes", "Bengali Faltu Memes", "Bengali Celebrity Memes", "Bengali Memes", "Bengali Funny Memes", "Bengali Political Memes", etc to search for these memes. To avoid copyright infringement, we only collect the memes from publicly accessible pages and groups. We accumulated a total of 7,532 memes from March 2022 to April 2023. The distribution of data sources is presented in Appendix B. We collected the memes with Bengali and code-mixed (Bengali + English) embedded texts. During the data collection, we filtered out memes if they (i) contained only visual or textual information, (ii) contained drawings or cartoons, and (iii) had uncleared contents either visually or textually. Some filtered meme samples are presented in Appendix Figure B.3. We also removed duplicate memes. After filtering and deduplication, we discarded 299 memes and ended up with a curated dataset of 7,233 memes. Following these, we use an OCR library (PyTesseract [^1] https://pypi.org/project/pytesseract/) to extract captions from the memes. As the captions have code-mixed texts, we manually reviewed and corrected if there were any missing words or spelling errors in the extracted captions. Finally, the memes and their captions are passed for manual annotation.
@@ 3.2 Dataset Annotation
We develop BHM focusing on two tasks: (i) detecting whether a meme is hateful or not and (ii) identifying the targeted entity of a hateful meme. We create guidelines clearly defining the tasks to ensure the annotation quality and mitigate the annotation bias. A few examples of annotated memes are depicted in Appendix Figure B.1.
@@ 3.2.2 Annotation Process
Two annotators independently annotated each meme, and the final label was determined through consensus between the annotators. On average an annotator spent 3 minutes to decide the label of a meme. In disagreement, an expert provided the final decision by discussing the uncertainties. For a minimal number of memes (< 2%), we observe that memes target multiple entities. Since this number is minimal for annotation simplicity, such samples were annotated with the dominant class label. During the final label assignment, we discarded 85 memes as the annotators and the expert could not agree on assigning a label. Finally, we get BHM, a multimodal Bengali hateful memes dataset with their targeted entities containing 7,148 memes.
@@
Inter-annotator Agreement: We computed the inter-annotator agreement in terms of Cohen’s Kappa ($\kappa$) score Cohen (1960) to check the validity. Table 1 shows the Kappa scores for each task category. We achieved a high agreement score of 0.79 for the hateful meme detection task. However, for target identification, we attained a moderate score of 0.63. These agreement scores indicate that annotators struggled to distinguish the targeted entities within hateful memes.
Figure 2: A simplified view of our proposed Dual Co-Attention Framework (DORA). The upper block represents the visual feature extractor and the lower block is the textual feature extractor. The Dual Co-Attention block takes encoded visual and textual representation and generates two attentive vectors namely VGAR (Vision-guided attentive Representation) and TGAR (Text-guided Attentive Representation). Finally, by concatenating the attentive vectors with the individual modality-specific features our method generates a richer multimodal representation.
@@ 3.3 Dataset Statistics
We divided the dataset into training (80%), validation (10%), and test (10%) sets for model training and evaluation. Table 2 shows the data distribution across different categories within each split. Task 1 exhibits a slight imbalance, while task 2 presents a significant imbalance, with most data falling under the ‘TI’ category. The distribution highlights that it will be challenging to accurately identify the targeted entities in hateful memes due to the limited number of samples in the ‘TO’, ‘TC’, and ‘TS’ categories. To acquire more insights into data characteristics, we analyze the training set memes. The analysis is presented in Appendix C.
@@ 4.1 Feature Extractor
To encode the visual information of the memes, we leverage the image encoder component of the CLIP (Contrastive Language Image Pretraining) Radford et al. (2021), a prominent visio-linguistic model. This image encoder incorporates a vision transformer Dosovitskiy et al. (2020) as its backbone. Meanwhile, we leverage the XGLM Lin et al. (2022), a multilingual generative language model, to encode meme captions. XGLM has demonstrated effectiveness in learning from diverse languages within context without parameter updates. Given that our dataset comprises code-mixed captions (Bangla + English), we posited that XGLM could offer a better-contextualized representation of these code-mixed captions. We fine-tuned the both image and text encoders, aiming to extract encoded representations. These representations were subsequently fed into a dual co-attention module to produce a multimodal representation.
@@ 4.2 Dual Co-Attention
To generate an effective multimodal representation, we fed the encoded visual and textual features into a dual co-attention block. Specifically, we generate two attentive multimodal feature representations using the Multi-head Self Attention (MSA) mechanism. The MSA takes three matrices: Query (Q), Key (K), and Value (V) as input. In standard NLP applications, all the matrices come from word representations. In contrast, motivated by Lu et al. (2019) we modified the MSA block where queries come from one modality and keys and values from another. This modification will generate an attention-pooled representation for one modality conditioned on another.
@@
To do this, we first weighed the visual features by performing a point-wise multiplication with attention scores and named it vision-guided attentive representation (VGAR). We also generate another attentive representation by weighing the textual features after observing the visual information. We called this text-guided attentive representation (TGAR). These two attentive representations (VGAR and TGAR) now contain significant cross-modal features. This cross-modal representation is further concatenated with the individual modality features (obtained from CLIP and XGLM). This process will boost the gradient flow and help the model learn from the individual as well as their cross-modal features. Finally, the combined multimodal representation is passed to the dense layer, followed by a softmax operation to predict the meme’s category.
@@ 5 Experiments
This section discusses the baselines and their performance comparison with the proposed method (DORA) and its variants. We developed several state-of-the-art computational models, including only visual models, only textual models, and multimodal models pre-trained on both modalities. We used weighted F1 score as the primary evaluation metrics. Other metrics such as precision and recall are also reported for the comparison. Appendix A provides the details of the experimental settings.
@@ 5.1.2 Multimodal Models
Table 3: Performance comparison of visual only, textual only, and multimodal models on the test set. Here, P, R, and F1 denote precision, recall, and weighted F1-score respectively. The VF, TF, VGAR, and TGAR denote the visual features, textual features, vision-guided attentive representation, and text-guided attentive representation respectively. The best-performing score in each column is highlighted in bold and the second-best score is underlined. The last row shows the performance improvement of the proposed system (DORA) over the best baseline score.

### v2

@@ Abstract
Internet memes have become a powerful means for individuals to express emotions, thoughts, and perspectives on social media. While often considered a source of humor and entertainment, memes can also disseminate hateful content targeting individuals or communities. Most existing research focuses on the negative aspects of memes in high-resource languages, overlooking the distinctive challenges associated with low-resource languages like Bengali (also known as Bangla). Furthermore, while previous work on Bengali memes has focused on detecting hateful memes, there has been no work on detecting their targeted entities. To bridge this gap and facilitate research in this arena, we introduce a novel multimodal dataset for Bengali, BHM (Bengali Hateful Memes). The dataset consists of 7,148 memes with Bengali as well as code-mixed captions, tailored for two tasks: (i) detecting hateful memes, and (ii) detecting the social entities they target (i.e., Individual, Organization, Community, and Society). To solve these tasks, we propose DORA (Dual cO-attention fRAmework), a multimodal deep neural network that systematically extracts the significant modality features from the memes and jointly evaluates them with the modality-specific features to understand the context better. Our experiments show that DORA is generalizable on other low-resource hateful meme datasets and outperforms several state-of-the-art rivaling baselines.
@@ 1 Introduction
In recent years, social media has brought a distinct form of multimodal entity: memes, providing a means to express ideas and emotions. Memes are compositions of images coupled with concise text. While memes are often amusing, they can also spread hate by incorporating socio-political elements. These hateful memes pose a significant threat to social harmony as they have the potential to harm individuals or specific groups based on factors like political beliefs, sexual orientation, or religious affiliations. Despite the significant influence of memes, their multifaceted nature and concealed semantics make them very hard to analyze. The prevalence of highly toxic memes in recent times has led to a growing body of research into the negative aspects of memes, such as hate Kiela et al. (2020), offensiveness Shang et al. (2021), and harm Pramanick et al. (2021b). However, most works focused on the memes in high-resource languages, while only a few studied the objectionable (i.e., hate, abuse) memes of low-resource languages Kumari et al. (2023). This is particularly true for Bengali.
@@
To bridge this research gap, we develop a novel Bengali meme dataset encompassing the targeted entities of hate. The captions in the dataset contain code-mixed (Bengali + English) and code-switched text (written Bengali dialects in English alphabets). This makes the dataset more distinctive and challenging compared to previous studies in the field. On the technical front, prior research on hateful meme detection Kiela et al. (2019); Pramanick et al. (2021a) revealed that off-the-shelf multimodal systems, which often perform well on a range of visual-linguistic tasks, struggle when applied to memes. Besides, most of the existing visio-linguistic models Radford et al. (2021); Li et al. (2019, 2022) are primarily trained on image-text pairs of English languages, thus limiting their capability on low-resource languages. Moreover, the existing state-of-the-art multimodal models Pramanick et al. (2021c); Lee et al. (2021) for hateful meme detection can not be replicated because several components of their architectures are not available in low-resource languages (i.e., Bengali). To tackle these issues, we developed a multimodal framework, and our major contributions are as follows:
  • We develop a benchmark multimodal dataset comprising 7,148 Bengali memes. The dataset includes two sets of labels for (i) detecting hateful memes and (ii) identifying targeted entities (individuals, organizations, communities, and society). We also provide detailed annotation guidelines to facilitate resource creation for other low-resource languages in this domain.
@@ 2 Related Work
Hateful memes dataset: Over the past few years, several meme datasets have been developed regarding hate speech and its various intensity levels, such as offense, harm, abuse, and troll. Sabat et al. (2019) developed a hateful memes dataset containing 5,020 memes. Facebook AI Kiela et al. (2020) contributed to developing a hateful memes dataset consisting of around 10K synthetic memes labeled in hateful and not-hateful classes. Similarly, another large-scale dataset comprising 150K memes was introduced by Gomez et al. (2020) for hate speech detection. In another work, Suryawanshi et al. (2020) developed an offensive memes dataset comprising 743 memes collected during the event of the 2016 US presidential election. Pramanick et al. (2021a) built a dataset to detect harmful memes containing around 3.5K related to COVID-19.
@@
Some works have also been conducted concerning low-resource languages. Perifanos and Goutsos (2021) developed a dataset of 4,004 memes for detecting hate speech in Greek. Kumari et al. (2023) developed an offensive memes dataset consisting of 7,417 Hindi memes. Recently, Das and Mukherjee (2023) introduced a dataset comprising 4,043 samples for detecting the Bengali abusive memes. Two prior studies focused on hateful meme detection in the Bengali language. Karim et al. (2022) introduced a synthetic hate speech dataset comprising around 4,500 Bengali memes. Similarly, Hossain et al. (2022) developed another Bengali hateful memes dataset having 4,158 memes labeled hateful or not-hateful. However, none of these datasets annotated the targets of hateful memes.
@@
Multimodal hateful meme detection: Over the years, various approaches have been employed for detecting hate speech using multimodal learning. Earlier researchers used conventional fusion (i.e., early and late) Suryawanshi et al. (2020); Gomez et al. (2020) produce producing multimodal representation. Later, some researchers employed bilinear pooling (Chandra et al., 2021) while others developed transformer-based multimodal architectures such as MMBT Kiela et al. (2019), ViLBERT Lu et al. (2019), and Visual-BERT Li et al. (2019). Besides, some works attempted to use disentangled learning (Lee et al., 2021), incorporate additional caption (Zhou et al., 2021), and add external knowledge Pramanick et al. (2021c) to improve the hateful memes detection performance. Recently, Cao et al. (2022) applied prompting techniques for hateful meme detection in English.
@@
Differences with existing studies: Though many studies have been conducted on hateful meme detection, only a few studies have focused on Bengali. We point out several drawbacks in the existing research on Bengali. Firstly, the existing hateful memes datasets are small. They framed the task as a binary (hateful or not-hateful) classification problem, overlooking the social entities (i.e., individuals, society) that a hateful meme can target. Our dataset provides both levels of annotation. The richness of our dataset contributes to a comprehensive understanding of the dynamics of Bengali memes. Only one particular work Das and Mukherjee (2023) studied the targets of abusive memes in Bengali. Their work is completely different from ours as we attempt to identify targeted social entities of hateful memes, which are more explicit than abuse. Secondly, the current research overlooked the memes containing cross-lingual captions, while many internet memes are written in code-mixed and code-switched manner. Lastly, none of the works provided any model that can be generalized across low-resource datasets.
@@ 3 BHM: A New Benchmark Dataset
As per our exploration, no dataset in Bengali currently focuses on capturing hateful memes that target specific entities. To fill this gap, we develop a novel multimodal hateful meme dataset. We followed the guidelines outlined by Hossain et al. (2022) and Kiela et al. (2020) to develop the dataset. This section will provide a detailed discussion of the dataset development steps, including the data collection and annotation process and relevant statistics.
@@ 3.1 Data Collection and Sampling
We collected memes from various online platforms, including Facebook, Instagram, Pinterest, and Blogs. We used keywords such as "Bengali Troll Memes", "Bengali Faltu Memes", "Bengali Celebrity Memes", "Bengali Memes", "Bengali Funny Memes", "Bengali Political Memes", etc to search for these memes. To avoid copyright infringement, we only collect the memes from publicly accessible pages and groups. We accumulated a total of 7,532 memes from March 2022 to April 2023. The distribution of data sources is presented in Appendix B. We collected the memes with Bengali and code-mixed (Bengali + English) embedded texts. During the data collection, we filtered out memes if they (i) contained only visual or textual information, (ii) contained drawings or cartoons, and (iii) had uncleared contents either visually or textually. Appendix Figure B.3 presents some filtered meme samples. We also removed duplicate memes. After filtering and deduplication, we discarded 299 memes and ended up with a curated dataset of 7,233 memes. Following these, we use an OCR library (PyTesseract [^1] https://pypi.org/project/pytesseract/) to extract captions from the memes. As the captions have code-mixed texts, we manually reviewed and corrected if there were any missing words or spelling errors in the extracted captions. Finally, the memes and their captions are passed for manual annotation.
@@ 3.2 Dataset Annotation
We develop BHM focusing on two tasks: (i) detecting whether a meme is hateful or not and (ii) identifying the targeted entity of a hateful meme. We create guidelines defining the tasks to ensure the annotation quality and mitigate the bias. Appendix Figure B.1 depicts a few annotated meme examples.
@@ 3.2.2 Annotation Process
Two annotators independently annotated each meme, and the final label was determined through consensus between the annotators. On average, an annotator spent 3 minutes deciding the label of a meme. In disagreement, an expert provided the final decision by discussing the uncertainties. For a minimal number of memes (< 2%), we observe that memes target multiple entities. Since this number is minimal for annotation simplicity, such samples were annotated with the dominant class label. During the final label assignment, we discarded 85 memes as the annotators, and the expert could not agree on assigning a label. Finally, we get BHM, a multimodal Bengali hateful memes dataset with their targeted entities containing 7,148 memes.
@@
Inter-annotator Agreement: We computed the inter-annotator agreement in terms of Cohen’s Kappa ($\kappa$) score Cohen (1960) to check the validity. Table 1 shows the Kappa scores for each task category. We achieved a high agreement score of 0.79 for the hateful meme detection task. However, for target identification, we attained a moderate score of 0.63. These agreement scores indicate that annotators struggled distinguishing the targeted entities within hateful memes.
Figure 2: A simplified view of our proposed Dual Co-Attention Framework (DORA). The upper block represents the visual feature extractor, and the lower block is the textual feature extractor. The Dual Co-Attention block takes encoded visual and textual representation and generates two attentive vectors: VGAR (Vision-guided attentive Representation) and TGAR (Text-guided Attentive Representation). Finally, our method generates a richer multimodal representation by concatenating the attentive vectors with the individual modality-specific features.
@@ 3.3 Dataset Statistics
We divided the dataset into training (80%), validation (10%), and test (10%) sets for model training and evaluation. Table 2 shows the data distribution across different categories within each split. Task 1 exhibits a slight imbalance, while task 2 presents a significant imbalance, with most data falling under the ‘TI’ category. The distribution highlights that it will be challenging to accurately identify the targeted entities in hateful memes due to the limited number of samples in the ‘TO’, ‘TC’, and ‘TS’ categories. We analyze the training set memes to acquire more insights into data characteristics. The analysis is presented in Appendix C.
@@ 4.1 Feature Extractor
To encode the visual information of the memes, we leverage the image encoder component of the CLIP (Contrastive Language Image Pretraining) Radford et al. (2021), a prominent visio-linguistic model. This image encoder incorporates a vision transformer Dosovitskiy et al. (2020) as its backbone. Meanwhile, we leverage the XGLM Lin et al. (2022), a multilingual generative language model, to encode meme captions. XGLM has effectively learned from diverse languages within context without parameter updates. Given that our dataset comprises code-mixed captions (Bangla + English), we posited that XGLM could offer a better contextualized representation of these code-mixed captions. We fine-tuned both the image and text encoders, aiming to extract encoded representations. These representations were subsequently fed into a dual co-attention module to produce a multimodal representation.
@@ 4.2 Dual Co-Attention
We fed the encoded visual and textual features into a dual co-attention block to generate an effective multimodal representation. Specifically, we generate two attentive multimodal feature representations using the Multi-head Self Attention (MSA) mechanism. The MSA takes three matrices: Query (Q), Key (K), and Value (V) as input. In standard NLP applications, all the matrices come from word representations. In contrast, motivated by Lu et al. (2019), we modified the MSA block where queries come from one modality and keys and values from another. This modification will generate an attention-pooled representation for one modality conditioned on another.
@@
To do this, we first weighed the visual features by performing a point-wise multiplication with attention scores and named it vision-guided attentive representation (VGAR). After observing the visual information, we also generate another attentive representation by weighing the textual features. We called this text-guided attentive representation (TGAR). These two attentive representations (VGAR and TGAR) now contain significant cross-modal features. This cross-modal representation is further concatenated with the individual modality features (obtained from CLIP and XGLM). This process will boost the gradient flow and help the model learn from the individual and their cross-modal features. Finally, the combined multimodal representation is passed to the dense layer, followed by a softmax operation to predict the meme’s category.
@@ 5 Experiments
This section discusses the baselines and their performance comparison with the proposed method (DORA) and its variants. We developed several state-of-the-art computational models, including only visual, textual, and multimodal models pre-trained on both modalities. We used weighted F1 scores as the primary evaluation metrics. Other metrics, such as precision and recall, are also reported for the comparison. Appendix A provides the details of the experimental settings.
@@ 5.1.2 Multimodal Models
Table 3: Performance comparison of visual only, textual only, and multimodal models on the test set. P, R, and F1 denote precision, recall, and weighted F1-score, respectively. The VF, TF, VGAR, and TGAR denote the visual features, textual features, vision-guided attentive representation, and text-guided attentive representation. The best-performing score in each column is highlighted in bold, and the second-best score is underlined. The last row shows the performance improvement of the proposed system (DORA) over the best baseline score.